### v2

# Improvement of Spiking Neural Network with Bit Planes and Color Models

Abstract
Spiking neural network (SNN) has emerged as a promising paradigm in computational neuroscience and artificial intelligence, offering advantages such as low energy consumption and small memory footprint. However, their practical adoption is constrained by several challenges, prominently among them being performance optimization. In this study, we present a novel approach to enhance the performance of SNN for images through a new coding method that exploits bit plane representation. Our proposed technique is designed to improve the accuracy of SNN without increasing model size. Also, we investigate the impacts of color models of the proposed coding process. Through extensive experimental validation, we demonstrate the effectiveness of our coding strategy in achieving performance gain across multiple datasets. To the best of our knowledge, this is the first research that considers bit planes and color models in the context of SNN. By leveraging the unique characteristics of bit planes, we hope to unlock new potentials in SNNs performance, potentially paving the way for more efficient and effective SNNs models in future researches and applications.
Index Terms:
Spiking neural network, spike coding, image classification
I Introduction
The field of artificial intelligence (AI) has experienced unprecedented advancements in recent decades[1], largely driven by the developments of deep learning algorithms. Despite these advancements, traditional artificial neural networks (ANNs) often fall short in modeling the intricate and dynamic nature of biological neural systems[2]. To bridge this gap, spiking neural networks (SNNs) have emerged as a promising alternative, leveraging the temporal dimension of neuronal activity to more closely emulate the behavior of biological brains.
SNNs utilize discrete spike events to transmit information between neurons, mirroring the asynchronous and event-driven communication observed in biological neurons[3]. This approach not only enhances the biological plausibility of neural models but also offers potential advantages in terms of computational efficiency and power consumption, particularly for hardware implementations[4]. The inherent sparsity and temporal coding of SNNs provide a robust framework for developing low-power neuromorphic systems that can operate efficiently in real-time environments [5].
Despite their potential, SNNs pose significant challenges, particularly in training methodologies and network optimization. Traditional backpropagation techniques used in ANNs are not directly applicable to SNNs due to their discontinuous and non-differentiable nature[6]. Consequently, novel training algorithms, such as spike-timing-dependent plasticity (STDP) [7, 8] and surrogate gradient methods [9], have been developed to address these challenges, enabling the effective training of SNNs.
Prior research in ANN has explored the application of bit planes (e.g. [10]) and color models [11, 12] to enhance performance across various tasks, demonstrating significant improvements in areas such as image recognition and processing. Motivated by these advancements, we aim to investigate whether bit planes and color models can similarly improve the performance of SNNs. This led us to formulate our research questions:
  • Does bit plane representation enhance the performance of SNN?
  • What are impacts of different color models on SNN?
Especially, we observe that bit planes of an image contain pulse-like information, and can be applied to encode color information for SNNs to possibly improve accuracy.// After extensive experimentation, we concluded that bit planes contain valuable information that can be utilized in the rate coding process as a supplementary source to enhance the accuracy of direct surrogate backpropagation process [13] in SNN architectures. Furthermore, bit plane information proved to be reusable across multiple color models in most of the cases examined in our study.
A preliminary version of this paper is published in [14], where the benefit of including bit plane coding is proved by three grayscale datasets. In this extended version, the following new points have been added. First, the proposed method was updated to support color images. Second, eight more datasets are included. Third, impacts of color models are investigated.
II Spiking Neural Network
In this section we present an overview of SNN knowledge related to our research.
II-A History of SNNs
Phenomenological neuron models can be progressively simplified until reaching a stage where even neural spikes are not modeled, leading to first and second generation neurons [15, 16]. This simplification significantly reduces computational demands but also decreases biological realism. The detail level in modeling often correlates with available computing power. For example, ANNs based on second generation neurons were historically limited by computing power [17], restricting the size of networks and the complexity of neuronal models. This limitation influenced the design of learning algorithms. Even with advancements in computing power, the complexity of neuronal models did not proportionately increase because existing learning algorithms were not compatible with more detailed models. Consequently, two distinct research fields emerged.
  • ANNs[18] focused on large networks of simplified neuron-like units, capable of learning through algorithms like backpropagation, and were successfully applied in pattern recognition, classification, and pattern completion tasks.
  • Computational Neuroscience[19] in contrast used detailed biophysical and phenomenological models in smaller networks to study electrophysiological processes, pattern generation, and dynamic behaviors of small neuron groups.
Although there have been attempts to create large networks of detailed biophysical neuron models [20], it remains challenging due to computational constraints.
Recent advances in computing power have blurred the lines between these fields. ANNs have become more biologically realistic, while networks of biophysical neurons have grown larger and more detailed. However, computational power still limits the application of detailed models in large neural networks for pattern recognition and classification. With improved computing power, new learning algorithms are being developed for these models. SNNs, which use spiking neurons modeled by phenomenological models like the Spike Response Model[21], represent the next step towards this goal. Due to the computational burden, biophysical models are less common in certain SNN applications. This paper focuses on research involving SNNs with phenomenological models, excluding those using biophysical models.
II-B The Integrate-and-Fire model
One of the earliest neuron models is the perfect integrate-and-fire (IF) model (also sometimes referred to as the non-leaky integrate-and-fire model), first studied by Louis Lapicque in 1907 [22]. In this model, the neuron is characterized by its membrane potential $V$:
$C\frac{dV(t)}{dt}=I(t)$ (1)
which evolves over time in response to an input current $I(t)$. This evolution is described by the time derivative of the capacitance law, $Q=CV$, where $C$ is the membrane capacitance. As an input current is applied, the membrane potential increases until it reaches a fixed threshold $V_{\text{th}}$. At this point, the neuron generates a spike modeled as a delta function, and the membrane potential is reset to its resting value, after which the process repeats. In this model, the firing rate increases linearly and indefinitely with the increase in input current.
While previous efforts have successfully converted ANN neurons with ReLU activations to IF neurons, challenges still remain in the conversion process. One of the key issues is the difficulty in precisely determining the appropriate thresholds in these conversion approaches. Despite the introduction of both data-driven and model-driven threshold determination methods[23], the converted SNN often experience unstable accuracy compared to their re-trained ANN counterparts. Additionally, IF neurons lack time-dependent memory. Once a neuron receives sub-threshold potentials, the membrane potential persists without decay until the neuron fires. This behavior is said to be inconsistent with biological neurons, which exhibit dynamic temporal properties [24]. Still, within certain SNN architectures, IF neurons have demonstrated superior performance compared to the supposedly better LIF neurons[24], with the SEW-ResNet[25] architecture being one such example.
Figure 1: Illustrated bit planes of a sample image from CIFAR10[26] dataset.
II-C Deep Residual Learning in Spiking Neural Networks
The success of ResNet[27] in deep learning has motivated efforts to extend residual learning to SNN. Prior attempts, commonly referred to as Spiking ResNets[28], have largely replicated the architecture of ResNet[27] by replacing ReLU activation functions with spiking neurons. However, this approach presents certain limitations:
  • Inability to achieve identity mapping across neuron models: Residual learning is predicated on the concept of identity mapping, where layers can be trained to approximate an identity function. While this is straightforward in ResNets[27] with ReLU activation, it is not always feasible in Spiking ResNets[28]. Specifically, [25] claimed that for certain spiking neuron models, such as leaky integrate-and-fire (LIF) neurons[22] with learnable membrane time constants, ensuring that a spiking neuron fires in response to an input spike becomes challenging. This complicates the realization of identity mapping.
  • Vanishing/exploding gradients: Even in cases where identity mapping can be achieved, Spiking ResNets[28] are prone to vanishing or exploding gradients during training. This occurs because the gradient propagation through multiple layers of spiking neurons may either diminish to zero or grow uncontrollably, complicating the training of deeper networks.
To address these issues, SEW ResNet[25] was introduced as a replacement design for residual blocks while utilizing IF neurons. The SEW block capitalizes on the binary nature of spikes by applying element-wise functions $g$ to the output of the residual mapping $A_{l}[t]$ and the input spike train $S_{l}[t]$[25]. This approach facilitates identity mapping, mitigates the vanishing/exploding gradient problem and achieves higher accuracy with fewer time steps compared to existing SNN models.
SEW ResNet [25] was also the first work to propose and apply a surrogate gradient approach for optimizing SNNs. Using backpropagation methods to train SNNs [29, 13] has become a common practice in neuromorphic computing. However, applying backpropagation in SNNs presents certain challenges. In SNNs, spike signals are typically modeled using binary values (0s and 1s) from a Heaviside function $H(\theta,t)$, where $\theta$ is the membrane potential and $t$ is a threshold:
$H(\theta,t)=\begin{cases}1&\text{if $\theta-t\geq 0$}\\ 0&\text{otherwise}\end{cases}$ (2)
The derivative $\partial H$ of this function is given by:
$\frac{\partial H}{\partial\theta}=\begin{cases}+\infty&\text{if $\theta-t=0$}\\ 0&\text{otherwise}\end{cases}$ (3)
As shown, directly differentiating $H(\theta,t)$ results in unstable gradients. To address this, [25] proposed substituting the derivative of the Heaviside function $\partial H$ with a numerically stable gradient from a different activation function during backpropagation, such as the derivative $\partial\sigma$ of the Sigmoid function $\sigma(\theta,t)$:
$\sigma(\theta,t)=\frac{1}{1+e^{\theta-t}}$ (4)
$\frac{\partial H}{\partial\theta}=\frac{\partial\sigma}{\partial\theta}=\frac{e^{\theta-t}}{(e^{\theta-t}+1)^{2}}$
This approach yields a more stable gradient during backpropagation, allowing the use of existing automatic differentiation frameworks to optimize SNNs effectively.
With these advantages, we decided to choose SEW Resnet[25] architecture for our study.
III Bit Planes and Color Models in Deep Learning
This section presents an overview of bit planes and color models, along with their applications in deep learning algorithms.
III-A Bit planes in deep learning
Bit planes have shown extensive utility in deep learning applications, offering new ways to improve model efficiency and performance. In [10], researchers used bit planes to encode the input layer in Binary Neural Network (BNN). They argued that bit planes retain crucial spatial information, enabling a reduction in the number of multiply-accumulate operations (MACs) without compromising the model’s accuracy or robustness. This demonstrates the potential of bit planes in streamlining computations while maintaining competitive performance.
The use of bit plane slicing on RGB images for adversarial robustness was explored in [30]. By selecting the most significant bit planes, the authors developed a classification model that was better equipped to handle adversarial attacks. This approach capitalizes on the spatial information encoded in different bit planes, which helps the model resist manipulations that would otherwise degrade its performance.
In the domain of medical imaging, bit plane slicing has also been applied to improve accuracy in breast cancer recognition tasks[31]. Researchers found that lower-order bit planes are more susceptible to noise, and by removing them before inputting the data into a convolutional neural network (CNN), they achieved higher classification accuracy. This method highlights how bit plane analysis can be tailored to specific challenges, such as noise reduction in medical images, to enhance model effectiveness.
Although bit planes have been widely studied and applied in ANNs, their utilization in SNNs is still unexplored. As far as we concern, our research is first to explore application of bit planes on SNNs spike coding process.
III-B Color models in deep learning
Color models are fundamental in computer graphics as they define how colors are represented, manipulated, and displayed for digital images. In computer graphics applications, understanding and effectively managing color models is crucial for tasks such as rendering realistic scenes, creating digital artwork, and designing user interfaces. Color models can be divided into device-oriented models, user-oriented models and device-independent models [32, 33].
Device-oriented color models are influenced by the signals of the devices and the resulting colors are affected by the technology used for display. Widely utilized in various applications that require color consistency with hardware tools, examples include any hardware devices used for human visual perception, such as televisions and video systems. Popular color models for devices include RGB, CMY(K) and YCbCr.
User-oriented color models are considered as a pathway between the observer and the device handling the color information. For our study, we mostly investigate cylindrically described models, including HSL and HSV. According to [32], these models also inherit other special properties such as device dependence and cylindrical-coordinate representation.
Device-independent color models are essential for defining color signals without reliance on the specific characteristics of any particular device or application. These models play a crucial role in applications that involve color comparisons and the transmission of visual information across networks that connect various hardware platforms. We will mostly research color models defined by the International Commission on Illumination (CIE), including CIE XYZ and CIE LAB.
Color models are becoming increasingly important in deep learning algorithms, offering valuable information that enhances model performance in various tasks. For example, in [34], a research on traffic light recognition tested a range of color models across different deep learning architectures, with results indicating that the RGB model remains favored despite some of its limitations. This suggests that, while RGB is not perfect, it still outperforms other models in certain applications.
In another study on applying convolutional neural networks (CNNs) for skin detection against adversarial attacks of generative models[35], RGB and YCbCr color models demonstrated superior performance, particularly in handling varying lighting conditions. This highlights the adaptability of these color models in scenarios where lighting inconsistencies can degrade model accuracy.
Moreover, the development of ColorNet[11] illustrates the potential of leveraging multiple color models to capture a broader set of features. By combining information from these diverse color models, ColorNet outperforms traditional models that rely on a single color model. However, this approach introduces computational overhead due to the inclusion of smaller submodels, which can be a challenge despite the model’s increased efficiency in some cases.
Despite extensive researches on the use of color models in ANN, their application in SNN remains underexplored and requires further investigation.
IV Proposed Method
In this section, we present in detail the design of our coding algorithm.
IV-A Hypothesis and Preliminary Study
We hypothesize that the information contained in the bit planes of an image can be effectively used to encode inputs for SNN training. Bit planes decompose an image into binary levels, each representing a specific bit of the binary representation of pixel values. This decomposition can potentially preserve critical information of the original image while providing a format that might be more suitable for the spike-based processing in SNNs. The granularity of information available in bit planes offers a unique approach to encoding visual data, which could enhance the learning capabilities of SNNs by providing a richer, multi-level representation of the input.
To test this hypothesis, we conducted a preliminary experiment and analysis using the MNIST[36], which is a simple dataset containing gray-scale images of hand-written numbers. For analysis in term of information contained in the coded spike signals, we perform the calculation of the mean base-2 Shannon entropy[37] $\overline{H}(X_{j})$ as:
$\overline{H}(X_{j})=-\frac{1}{N}\sum_{j=1}^{N}\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (5)
and the approximated average Kolmogorov complexity $\overline{K}$ across all dataset’s elements where the Kolmogorov complexity $K(X)$ of a coded image is estimated with DEFLATE compression algorithm[38] $C_{DEFLATE}(X)$ as:
$\overline{K}(X_{j})$ $=\frac{1}{N}\sum_{j=1}^{N}K(X_{j})$ (6)
$\approx\frac{1}{N}\sum_{j=1}^{N}C_{DEFLATE}(X_{j})$
$=\frac{1}{N}\sum_{j=1}^{N}Huffman(LZ77(X_{j}))$
where $Huffman(X)$ is the Huffman coding[39] and $LZ77(X)$ is the LZ77 compression[40] (or sometimes refered to as sliding window compression).
We compared three approaches: traditional spike coding, bit plane coding only, and the combination of both spike coding and bit plane coding. The detailed settings of this experiment can be found in Section V-A. The results are provided in Table I.
Coding method Validation Accuracy (%) Averaged Kolmogorov complexity approximation (bits) Averaged Shannon entropy (bits)
Spike coding 98.62 721.5 0.53
Bit plane coding only 97.87 772.8 0.56
Combination of spike coding and bit plane coding 98.93 744.4 0.55
TABLE I: Average validation accuracy and analysis metrics comparison on MNIST[36]
The results demonstrate that the approximated Kolmogorov complexity and Shannon entropy of bit-plane-coded data are higher, suggesting that the information content is more complex and less predictable. This may help explain why the accuracy of models using exclusively bit-plane-encoded inputs is lower compared to the baseline. On the other hand, a combination of rate coding and bit-plane coding resulted in an increase in performance. This outcome demonstrates that bit planes contain valuable information that SNN can learn from to improve their classification ability. This encourages us to carry out a further study and perform testing on other datasets.
IV-B Proposed Coding Method with Bit Planes
Figure 2: Our proposed coding method (lower) compare with traditional spike coding method (upper).
As we can see in Figure 2, the proposed coding scheme integrates a dual approach featuring both traditional spike coding and bit plane coding methodologies. Initially, the process commences with the conventional spike coding procedure applied to the image tensor. The bit plane coding algorithm (Algorithm 1) is employed to transform an input image into binary planes. The final encoded tensors are then generated through tensor concatenation across the first dimension, merging the outputs from spike coding and bit plane coding. These encoded tensors are thereby prepared for utilization in inference tasks and/or training processes.
Algorithm 1 is designed to encode an image tensor batch $X$ into $n_{bit}$ bit planes. To determine $n_{bit}$, it is essential to know the maximum possible value $X_{max}$ of the tensor batch. Since different color models have different $X_{max}$, this value must be predefined according to the selected color model. Once $X_{max}$ is defined, $n_{bit}$ can be calculate as $\left\lceil\log_{2}(X_{max})\right\rceil$.
To calculate the bit planes of an unnormalized integer input tensor batch $X$, we perform element-wise division of $X$ by 2. The remainder from this division constitutes a bit plane of $X$, while the quotient is saved for the extraction of subsequent bit planes. This process is repeated until we obtain $n_{bit}$ bit planes. The resulting bit planes are then concatenated to form $n_{bit}$ binary tensor batch of $X_{shape}$. This format closely resembles that of rate coding and ensures determinism in the coding results without having to sample more than $n_{bit}$ times.
Symbol Meaning
$X$ Input image tensors
$X_{max}$ Integer value indicate the maximum value of input image tensor
$n_{bit}$ Minimum number of bit planes required to encode input $X$ into bit planes
$log_{2}()$ Binary logarithm function
$X_{shape}$ Dimensions of X, usually include number of images being processed $B$, number of channels $C$ and resolution $H,W$
$concat()$ Tensor concatenation function, concatenating a list of tensors resulted to a tensor contain spikes of shape ($n_{bit}$, $X_{shape}$)
$L$ A list-like [41] data structure, used to temporarily store coding result
$\mathbb{Z}$ The set of all integers
TABLE II: Table of symbols used in Algorithm 1.
Algorithm 1 Bit plane coding algorithm
$X_{max}$ is defined, $X_{max}>0$ and $X\in\mathbb{Z}$
function BitplaneEncode($X,X_{max}$)
     $n_{bit}\leftarrow\left\lceil\log_{2}(X_{max})\right\rceil$triangleright Get number of bit required to encode image from highest possible value $X$
     $L\leftarrow[\text{ }]$triangleright Empty list initialization
     for each $X_{shape}.C$ do
         for $i\leftarrow 0,i<n_{bit},i++$ do
              $L.insert(X\mod{2})$triangleright Element-wise modulo of tensor, then store the bit plane to $L$
              $X\leftarrow\left\lfloor X/2\right\rfloor$triangleright Element-wise division of input image tensor and floor the result
         end for
     end for
     return $concat(L)$triangleright Concatenate all stored sub tensors in $L$ to a single tensor
end function
V Experiments
V-A Experimental settings
To investigate performance of SNN for images, we employed the popular tool SEW-ResNet18 with ADD element-wise function [25]. The model and experiment process were implemented using PyTorch [42] and SpikingJelly[43]. Color conversion algorithms are based on implementations from Kornia [44]. For the baseline method in the experiments, the spike rate coding is employed with a step size of 10. To facilitate optimization, we utilize surrogate gradient methods applied to all IF neurons using the arctan surrogate function $S(U)$ [25, 13] with the defined derivative $\partial S$ of:
$\frac{\partial S}{\partial U}=\frac{1}{\pi}\cdot\frac{1}{1+(\pi U)^{2}}$ (7)
The optimization process is conducted with the Adam algorithm [45], using a learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ across all evaluated datasets. Cross-Entropy Loss with Softmax function and mean reduction $L(x,y)$[42] across $N$ mini-batch dimension was also widely employed as the loss function:
$L(x,y)=\frac{\sum_{n=1}^{N}l_{n}}{N},$ (8)
$l_{n}=-\sum_{c=1}^{C}w_{c}\log\frac{\exp{x_{n,c}}}{\sum_{i=1}^{C}\exp{x_{n,i}}}y_{n,c}$
where $l_{n}$ is loss of each mini-batch, $x$ is the input, $y$ is the target, $w$ is the weight, $C$ is the total amount of class where $i\in C$ and $c$ is the mini-classes.
In the first part of our experiments, we only use grayscale image datasets to clearly see the effects of bit plane coding. In the second part, popular datasets of color images are used to see the practical effects of both bit plane coding and color models. We allocated 80% of each dataset for training and the remaining 20% with deterministic random seeding for validation. All datasets were processed with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU.
V-B Experimental Results
In this part, we will discuss results of our main experiments along with some interesting findings beside original hypotheses.
V-B1 The effect of bit plane coding on grayscale images
Dataset Accuracy of coding methods (%)
Baseline Bit plane coding only Proposed method
MNIST[36] 98.62 97.87 98.93
KMNIST[46] 94.13 92.33 95.58
Fashion-MNIST[47] 88.83 86.66 90.19
TABLE III: Comparison of validation accuracy among three coding methods. Best value is highlighted in bold.
Dataset Analysis results of coding methods (bits)
Baseline Bit plane coding only Proposed method
AKC SE AKC SE AKC SE
MNIST[36] 721.5 0.53 772.8 0.56 744.4 0.55
KMNIST[46] 1093 0.70 1221 0.70 1150 0.70
Fashion-MNIST[47] 1384 0.86 1377 0.82 1381 0.84
TABLE IV: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE). Highest values of AKC are denoted in red and lowest are denoted in blue.
First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information. For that, we employed three popular datasets of grayscale images, including MNIST[36], KMNIST[46], Fashion-MNIST[47]. The results of MNIST dataset were already shown in Table I. However, they are still included here to see the behaviors of the coding methods over different datasets. As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel. This process would produce eight bit-plane channels that represent the image information in binary form. In this part, we investigate three options, namely 1) the baseline model (spike coding only), 2) using bit plane coding only, 3) the proposed model (spike coding combined with bit plane coding).
As demonstrated in Table III, the combination of spike coding with bit-plane coding outperformed other coding methods in terms of validation accuracy. As mentioned, using only bit-plane coding resulted in a reduction in accuracy; however, the combination of spike coding with bit plane coding (i.e. the proposed method) provides gain in all datasets. For any method, the accuracy value is highest with the MNIST dataset and lowest with the Fashion-MNIST dataset. This could be because among these three datasets, MNIST is the simplest one while Fashion-MNIST is the most complex one. Interestingly, the gain of the proposed method compared to the baseline is increased from 0.31% (with MNIST) to roughly 1.4% (with KMNIST and Fashion-MNIST). That means the impact of bit plane coding is more prominent with more complex datasets.
Upon further analysis of spike distribution complexities for various coding methods, as presented in Table IV, several insights emerge.
First, we observe that the baseline method has the lowest complexity in MNIST and KMNIST datasets, while it has the highest complexity in the Fashion-MNIST dataset. We suspect that this difference stems from the nature of these datasets. Both MNIST and KMNIST primarily use binary pixel values, with zeros representing the background and ones indicating the presence of a character. This binary representation results in a simpler spike distribution. In contrast, Fashion-MNIST employs a wider range of pixel values to capture the textures of various clothing items, which introduces greater complexity into the spike distribution when using baseline coding methods. This observation suggests that more complex visual datasets may lead to even more pronounced differences in spike distribution when using these coding methods.
Additionally, by introducing a hybrid coding approach that combines baseline and bit plane coding methods, we hypothesize the potential existence of an optimal complexity threshold at which the SNN might achieve peak performance. This idea of an optimizable global maximum suggests that complexity thresholds could play a key role in determining SNN effectiveness. Based on our data, we propose that measuring spike distribution complexity could be crucial in tuning SNNs to perform optimally across different datasets and tasks.
V-B2 The effect of bit plane coding on color images
Dataset Accuracy of coding methods (%)
Baseline Bit plane coding only Proposed method
CIFAR10[26] 70.69 37.28 73.49
CIFAR100[26] 38.57 11.28 42.15
Caltech101[48] 61.67 43.28 64.55
Caltech256[49] 30.04 21.75 41.51
EuroSAT[50] 84.39 79.87 88.48
Imagenette[51] 72.99 48.20 78.68
Food101[52] 19.62 6.29 36.95
GTSRB[53] 93.36 95.85 95.35
TABLE V: Average validation accuracy on various computer vision dataset of our proposed method.
Dataset Analysis results of coding methods (bits)
Baseline Bit plane coding only Proposed method
AKC SE AKC SE AKC SE
CIFAR10[26] 7530 0.995 6764 0.996 7190 0.995
CIFAR100[26] 7374 0.995 6710 0.996 7079 0.995
Caltech101[48] 310k 0.993 260k 0.992 279k 0.992
Caltech256[49] 307k 0.991 257k 0.990 276k 0.990
EuroSAT[50] 30.3k 0.953 22.9k 0.944 27.0k 0.949
Imagenette[51] 334k 0.987 282k 0.994 302k 0.992
Food101[52] 339k 0.989 288k 0.995 308k 0.993
GTSRB[53] 308k 0.902 2914 0.223 5791 0.019
TABLE VI: Comparison of analysis results among coding methods, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE), values in thousands are represented with k (e.g., 22k = 22,000). Highest values of AKC are denoted in red and lowest are denoted in blue.
Dataset Analysis results of coding methods (bits)
With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE AKC SE
CIFAR10[26] 7190 0.995 7086 0.992 6671 0.972 6634 0.927 6752 0.957 7083 0.986 6857 0.988
CIFAR100[26] 7079 0.995 6975 0.991 6606 0.974 6551 0.929 6660 0.955 6984 0.987 6769 0.985
Caltech101[48] 279k 0.992 269k 0.978 231k 0.957 241k 0.905 242k 0.929 268k 0.989 239k 0.971
Caltech256[49] 276k 0.990 265k 0.976 233k 0.957 243k 0.907 244k 0.933 244k 0.989 242k 0.972
EuroSAT[50] 27.0k 0.949 26.5k 0.968 25.2k 0.944 25.3k 0.993 25.8k 0.950 26.2k 0.932 25.9k 0.966
Imagenette[51] 302k 0.992 292k 0.989 267k 0.957 271k 0.901 276k 0.941 288k 0.979 277k 0.976
Food101[52] 308k 0.993 297k 0.990 273k 0.986 277k 0.908 280k 0.927 294k 0.975 281k 0.969
GTSRB[53] 5791 0.019 5706 0.006 17.6k 0.489 86.1k 0.462 92.9k 0.499 5704 0.005 5703 0.006
TABLE VII: Comparison of analysis results with our proposed coding method across all tested color models, contains the approximated Kolmogorov complexity (abbreviated AKC) and Shannon Entropy (abbreviated SE) measured in bits, values in thousands are represented with k (e.g., 22k = 22,000). Highest values of AKC are denoted in red and lowest are denoted in blue.
Dataset Accuracy of coding methods (%)
With RGB encoded (Ours) With CMY encoded (Ours) With YCbCr encoded (Ours) With HSL encoded (Ours) With HSV encoded (Ours) With CIE XYZ encoded (Ours) With CIE LAB encoded (Ours)
CIFAR10[26] 73.49 73.73 73.64 73.39 73.06 73.63 74.06
CIFAR100[26] 42.15 41.31 41.62 41.94 41.91 41.73 42.03
Caltech101[48] 64.55 65.65 65.53 66.46 68.41 65.30 66.22
Caltech256[49] 41.51 40.65 40.84 40.55 39.94 40.50 40.79
EuroSAT[50] 88.48 86.61 88.11 88.74 88.81 88.87 87.48
Imagenette[51] 78.68 77.35 78.78 78.55 78.50 77.35 78.70
Food101[52] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
GTSRB[53] 95.35 51.36 70.68 79.18 80.86 51.78 51.78
Average 65.15 59.21 62.02 63.36 63.49 59.53 59.61
TABLE VIII: Comparison of validation accuracy among coding methods.
For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[26], CIFAR100[26], EuroSAT[50], Caltech101[48], Caltech256[49], Imagenette[51] (which is a subset of ImageNet[54]), GTSRB[53], and Food101[52]. The default color model of these datasets is RGB. Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels. So, among these datasets, CIFAR10[26], CIFAR100[26] and EuroSAT[50] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
Across all tested datasets, our proposed methods consistently outperform the baseline method, with the notable exception of the GTSRB[53] dataset, where bit-plane coding alone appears to yield the best performance without the need for combination with traditional spike coding. This observation suggests that while our methods generally enhance performance by improving the representation of spike distributions, certain datasets may favor different coding approaches depending on their visual complexity and specific characteristics. In the GTSRB[53] case, it seems that bit-plane coding alone is sufficient to capture the essential features for effective processing.
The analysis results presented in Table VI highlight the relationship between complexity and performance in different coding methods. Bit-plane coding consistently exhibits lower complexities across datasets, while traditional spike coding tends to produce higher complexity values. This outcome aligns with our hypothesis that complexity increases in visual datasets with greater detail and variation.
In further investigating the GTSRB[53] dataset, we observed a significant reduction in spike distribution complexity (over 99% when comparing baseline with proposed scheme and pure bit planes coding in Table VI) alongside improved model performance. Usually a decrease in complexity can sometimes indicate a loss of critical information, it may also suggest that the spike distribution has become more streamlined, capturing essential features more effectively. However, a complexity reduction at such large scale could also introduce elements of instability (shown in Table VII), as it can be challenging to discern whether the decreased complexity represents a beneficial simplification or a detrimental loss of information.
Consequently, while our findings indicate promising results, we suggest further exploration of complexity thresholds to ensure consistent performance across varying datasets.
V-B3 Impacts of color models
In this part, also we investigate two options, namely 1) the baseline method and 2) proposed method.
The accuracy results of different cases are shown in Table VIII. We can see that, when using RGB color model, the accuracy of the proposed method is always higher than that of the baseline. The gain ranges from about 2% (with GTSRB[53] dataset) up to 11% (with Caltech256[49] dataset), depending on the employed dataset. This suggests that the benefit of the proposed method when applying to color images is more significant than when applying to grayscale images.
Interestingly, the results show that RGB is not always the best color model. Among the eight datasets, RGB provides the best performance in three cases (namely CIFAR100[26], Caltech256[49], GTSRB[53]). CMY does not achieve any best results, whereas each of the other five color models achieves the best performance once. Especially, with Caltech101[48] dataset, the accuracy of HSV color model is 68.41%, which is much higher than that of RGB color model (64.55%).
When considering the average of accuracy values across all datasets, the RGB color model is significantly better than other color models (mentioned the last row of Table VIII) and even though there are performance degradation on GTSRB[53], averaged result across all dataset proven to at least out performed baseline method.
From the above discussion, the key findings of our research can be summarized as follows.
  • The proposed method provides benefits for both grayscale images and color images.
  • The gain of the proposed method depends on the employed dataset. For practical color datasets, the gain is from 2% up to 11%.
  • In terms of average accuracy value over all datasets, RGB is the best color model and CMY is the worst one.
  • However, considering individual datasets, RGB is not always the best color model. It provides the best performance in only three datasets.
  • Especially, with the Caltech101 dataset, the performance of HSV color model (68.41%) is much better than that of the RGB color model (64.55%).
VI Discussion
VI-A Complexity or entropy?
In this study, we employ two primary metrics, an approximated Kolmogorov complexity and Shannon entropy, to analyze the effects of bit-plane coding on the distribution of spike trains. These metrics were chosen because they are widely used in signal processing and information theory, making them well-suited for analyzing the coded spike and bit-plane signals. By treating both coded spikes and bit planes as types of signals, we believe that Kolmogorov complexity and Shannon entropy can provide valuable insights into the complexity of spike distributions. While we conducted limited analysis with Shannon entropy, our results indicate a positive correlation between Shannon entropy and Kolmogorov complexity. This suggests that Shannon entropy could serve as an alternative to Kolmogorov complexity for those interested in measurement and analysis. However, we find Kolmogorov complexity to be more effective, as it provides clearer distinctions between measurements when compared to Shannon entropy.
Despite achieving substantial insights through these metrics, we recognize that measuring complexity and entropy alone may not fully capture the intricacies of image data. Images inherently contain structural complexity and spatial information, which can be difficult to quantify and analyze. Focusing solely on spike signals as sequential data may overlook these aspects. A helpful analogy is the historical debate in physics regarding the nature of light, with some scientists viewing light as a wave and others as particles. It was only with the discovery of wave-particle duality through experiments like the double-slit experiment that a more comprehensive understanding emerged. Similarly, viewing spike signals solely as sequential signals could reveal certain patterns but may not capture the full complexity underlying the optimization process in image processing tasks. Currently, tools designed specifically for spike signal analysis that incorporate spatial information for image-based tasks are underexplored. This gap highlights the need for further research into analysis methods that can integrate both signal and spatial information for more complete insights.
VII Conclusion
In conclusion, our research introduces a novel coding method for SNN that leverages bit planes derived from various color models of input image data. This approach prove to enhances the computational accuracy of SNN without increasing the model size. Through extensive experimental validation, we have demonstrated the effectiveness of our coding strategy across multiple computer vision tasks, achieving notable performance gains compared to conventional methods. This study is pioneering in its application of color models within the context of spiking neural networks, and we believe that our findings open new avenues for the development of more efficient and effective SNN models. Our work sets a foundation for future research and applications, potentially advancing the field of computational neuroscience and artificial intelligence.
Appendix A CMY or CMYK?
In the initial stages of our experimental setup and algorithm design, we encountered challenges in selecting the appropriate device-dependent color model for our experiments. Specifically, we debated whether to use the CMY or CMYK color model. Each has advantages in different applications, but their compatibility with the neural networks used in our research required careful consideration. In image processing and printing, the CMYK model was developed to improve print quality by reducing ink usage in darker regions. This is achieved by adding a K component, which reduces the CMY values proportionally, leading to enhanced contrast and more economical ink usage in printed images.
However, when adapting this approach for digital image processing, particularly for direct usage on digital devices, the inclusion of a K channel becomes less beneficial. Unlike printed images, digital displays do not require ink density adjustments, making the K channel unnecessary for accuracy. For our experiment, this would introduced dimensional conflicts in tested SNN architectures and coding schemes since they are previously optimized for three-channel data inputs (RGB), including a fourth channel would require additional modifications to model architecture, potentially affecting performance and comparability.
Furthermore, in our image processing pipeline, we observed that converting images to the CMYK color model using standard image processing libraries (e.g., PIL[55]) did not apply undercolor removal during conversion to CMYK model, resulting in a K channel of zero values. This posed a unique problem in our training setup: a zeroed K channel in the input data could nullify both the input signal and its gradients when surrogate gradients are used for backpropagation in SNNs. Consequently, this would undermine the learning process by reducing the effectiveness of gradient-based updates, complicating model training and ultimately impacting the accuracy of our results.
Appendix B Tranning progession details
This section contain the training progression graph for each variant of experiments in Section V-B3 and Section V-B1.
It is noteworthy from Figure 4 and 5 that performance fluctuations were observed in both traditional coding and our coding method. We hypothesize that these fluctuations might related to inherent error rate during sampling process in Poisson-distributed spike trains[3].
Another hypothesis is that although bit planes may possess a certain degree of spatial information, they can also encompass noise on low order bit planes that interferes with the training process or have little to none distinguishable features (which might explain how on low noises single-channel image dataset such as MNIST, the performance of only using bit planes coding remain competitive to traditional rate coding with less amount of time step). Consequently, the selective identification of bit planes that retain valuable spatial information continues to pose a significant challenge.
(a) MNIST
(b) Fashion-MNIST
(c) KMNIST
Figure 3: Comparison in accuracy progression and loss progression between coding methods on different low-resolution gray-scale image classification datasets.
(a) CIFAR10[26]
(b) CIFAR100[26]
(c) Caltech101[48]
(d) Caltech256[49]
(e) EuroSAT[50]
(f) Imagenette[51]
(g) Food101[52]
(h) GTSRB[53]
Figure 4: Comparison in loss progression of traditional rate coding with our method on image classification datasets. On every sub-figure, loss details of every color models is shown on the left and average result of all models on the right with maximum loss difference denoted on top of figure.
(a) CIFAR10[26]
(b) CIFAR100[26]
(c) Caltech101[48]
(d) Caltech256[49]
(e) EuroSAT[50]
(f) Imagenette[51]
(g) Food101[52]
(h) GTSRB[53]
Figure 5: Comparison in accuracy progression of traditional rate coding with our method on image classification datasets. On every sub-figure, accuracy details of every color models is shown on the left and average result of all models on the right with maximum accuracy difference denoted on top of figure.